Find (–7) – 8 – (–25) Find (–7) – 8 – (–25) Find (–7) – 8 – (–25)

 Find 
(–7) – 8 – (–25) 
Find 
(–7) – 8 – (–25) 
Find 
(–7) – 8 – (–25) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(-7)–8–(-25) = -7 -8 + 25
= -15 + 25
= 10

Thanks